OpenStudy (anonymous):

Find the slope of the line 2x+4y=-1 .

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

i got the same thing too...

OpenStudy (anonymous):

2x+4y=-1 . --- 4y= -2x-1 ===> y=(-1/2)x -1/4 ===> the slope = -1/2